Given the point (2,-4) and a slope of -1/2 find the equation of a line having this slope and point.
Answer by rmromero(383)

To find the equation of the line given a point and slope substitute the given informations

Let (x, y) = (2, -4) 
         m = -1/2

           y - y1   =   m (x - x1)
           y - (-4) = -1/2 (x - 2)
             y + 4  = -1/2x + 1
                 y  = -1/2x - 3 <----- Equation written in slope-intercept form
                2y  = -x - 6
         x + 2y + 6 = 0      <------- Equation written in standard form